how does charge redistribute when a charged parallel plate capacitor is connected to another identical capacitor in series?

Grade:12

1 Answers

Khimraj
6 years ago
Let the charge on first capactor of capacitance C is Q.
Since capacitor are connected in series.
V = V1 + V2 = Q1/C1 + Q2/C2
here C1 = C2 =C
Q1 = Q and Q2 = 0
So V = Q/C
So charge on each capacitor will be C*(V/2) = Q/2.
So total charge remains constant but iy will equally distribute among both.
Hope it clears. If you like answer then please approve it.